According to Chebyshev's Theorem, at least what proportion of the data will be within 1.2 standard deviation distance about the mean? 27% 30% 61% 68% 69%
Answer #1

Answer

we know the formula

Area within k standard deviation =1 - 1/k^2

we have to find area or proportion within 1.2, so putting k = 1.2

we get

proportion = 1 - 1/1.2^2

= 1 - 0.6944

= 0.3056

= 0.3056*100

= 30.56%

= 30% (approx)

option B is correct (30%)